OpenStudy (anonymous):

Please help!! Write the rate expressions for the following reactant in terms of disappearance of the reactants and the appearance of the products C6H12O6 + 6O2 --> 6CO2 + 6H2O

OpenStudy (aaronq):

you can write them as differentials, you only need to indicate the coefficients and the sign - negative for disappearance. \(\sf rate=\pm \dfrac{1}{coefficient}*\dfrac{d[A]}{dt}\) for example, for oxygen, \(\sf overall~rate=-\dfrac{1}{6}*\dfrac{d[O_2]}{dt}\)